Durban Moving and Storage wants to have enough money available 6 years from now to purchase a new tractor-trailer. If the estimated cost will be $210,000, how much should the company set aside each year if the funds earn 10% per year? The company should set aside $ 35000 35000 Incorrect each year.
Answer #1

The formula here to be used is Recurring deposit formula = P*(((1+r)^n)-1)/(1-(1+r)^(-1))

P is regular deposit

r is annual interest rate

n is no of years

210000=P*(((1+0.1)^6)-1)/(1-(1+0.1)^(-1/1)) = P*8.49

210000/8.487 = 24743.73

So, P = 24743.73 needs to be set aside every year